OpenStudy (anonymous):

Guys, i need an instant answer. If i have a function y=(b*x-a)/(1+c*x), how can i find x when i already have y? For example, y=33.6, how much is x?

RhondaSommer (rhondasommer):

hmmm hold on

RhondaSommer (rhondasommer):

write an actual equation using the little equation button below or take a pucture of it so I know what you are doing. that way i know if i can help you or not

OpenStudy (anonymous):

I written an actual equation

RhondaSommer (rhondasommer):

that was an example of an actual equation.

RhondaSommer (rhondasommer):

\[y=\frac{(b^x-a)}{(1+c^x)}\] is it that?

OpenStudy (anonymous):

No

RhondaSommer (rhondasommer):

\[y=\frac{b^{x-a}}{(1+c*x)}\]

RhondaSommer (rhondasommer):

\[y=\frac{(b*x-a)}{(1+c*x)}\] or that?

OpenStudy (anonymous):

Yes

RhondaSommer (rhondasommer):

then you cannot solve it because you need (b) and (a). without those two solving the problem mathmatically is impossible.

RhondaSommer (rhondasommer):

you would have ≠ no answer.

OpenStudy (anonymous):

i have them

RhondaSommer (rhondasommer):

wow. now you tell me. okay tell me what they are then.

OpenStudy (anonymous):

B= 0,187344 and a= -3,5169955 and c= -0,014443694

RhondaSommer (rhondasommer):

and your y is 33.6?

OpenStudy (anonymous):

yes

RhondaSommer (rhondasommer):

so \[33.6=\frac{\left(0.18734x+3.517\right)}{\left(1-0.01444x\right)}\]

RhondaSommer (rhondasommer):

\[\mathrm{Multiply\:both\:sides\:by\:}1-0.01444x\]

RhondaSommer (rhondasommer):

\[33.6\left(1-0.01444x\right)=\frac{0.18734x+3.517}{1-0.01444x}\left(1-0.01444x\right)\]so then it is \[33.6\left(1-0.01444x\right)=0.18734x+3.517\]

RhondaSommer (rhondasommer):

\[distribute~33.6\left(1-0.01444x\right)\]\[33.6\left(1-0.01444x\right)=0.18734x+3.517\] \[\mathrm{Multiply\:both\:sides\:by\:}1000000\]\[33.6\cdot \:1000000-0.48518x\cdot \:1000000=0.18734x\cdot \:1000000+3.517\cdot \:1000000\]\[33600000-485184x=187340x+3517000\]

RhondaSommer (rhondasommer):

\[\mathrm{Subtract\:}33600000\mathrm{\:from\:both\:sides}\]\[33600000-485184x-33600000=187340x+3517000-33600000\]\[-485184x=187340x-30083000\] \[\mathrm{Subtract\:}187340x\mathrm{\:from\:both\:sides}\]\[-485184x-187340x=187340x-30083000-187340x\]\[-672524x=-30083000\] \[\mathrm{Divide\:both\:sides\:by\:}-672524\]\[\frac{-672524x}{-672524}=\frac{-30083000}{-672524}\]\[x=\frac{7520750}{168131}\]\[\mathrm{Show\:solutions\:in\:decimal\:form}\]\[x=44.73149\]

RhondaSommer (rhondasommer):

there ya go.

OpenStudy (anonymous):

but its not correct. It has to be somewhere between 50 and 60 definetly

RhondaSommer (rhondasommer):

well its not. soooo

OpenStudy (anonymous):

Looking at graph, it is 54

OpenStudy (anonymous):

excuse me, 55

OpenStudy (anonymous):

Could you write me a solution with just a b c and x without numbers. I would be very gratefull.

RhondaSommer (rhondasommer):

http://www.mathpapa.com/algebra-calculator.html look

RhondaSommer (rhondasommer):

you chose 33.6 dont forget that.

OpenStudy (anonymous):

haha thank you for this website...

RhondaSommer (rhondasommer):

your welcome. but it was more to prove that my calculations were correct

OpenStudy (anonymous):

But i still dont get it.

RhondaSommer (rhondasommer):

either you chose the wrong y, or have the wrong equation

OpenStudy (anonymous):

for 45% i have 15, and for 33.6 i have now 44%. Its not possible

OpenStudy (anonymous):

This equation works, as i found y to be correct with any x

OpenStudy (anonymous):

Ill put 44 and find out y value

RhondaSommer (rhondasommer):

but is it? here let me do one with x and see if it is really correct

OpenStudy (anonymous):

If you put x to be 44, the y is 12.96

RhondaSommer (rhondasommer):

i got y=32.25088 for x=44

RhondaSommer (rhondasommer):

its your equation mate :/

OpenStudy (anonymous):

x=44 cant be 32.... i did it 5 times

OpenStudy (anonymous):

The problem is that that site is wrong. It gives me a solution for x, but clearly it isnt

OpenStudy (anonymous):

this is so strange

OpenStudy (anonymous):

If i start to guess x and plug in the formula, i will eventualy get close value of y, but there just HAS to be a better way

RhondaSommer (rhondasommer):

my bad wrong link http://www.cymath.com/answer.php?q=y%3D(0.18734(44)%2B3.517)%2F(1-0.01444(44))

OpenStudy (anonymous):

I tried for x=55 and i got y=33. So the formula corresponds to the graph. Everything is fine exept the equation for x

RhondaSommer (rhondasommer):

my bad wrong link http://www.cymath.com/answer.php?q=y%3D(0.18734(44)%2B3.517)%2F(1-0.01444(44))

OpenStudy (anonymous):

Ok, but that link contains the equation for y, but not for x

RhondaSommer (rhondasommer):

jsyk i rounded x and y :)

OpenStudy (anonymous):

do you know where is wrong?

RhondaSommer (rhondasommer):

up to 32 and 44. when x is 44 y is (about) 32 when y is 32 x is (about) 44

OpenStudy (anonymous):

But ask yourself, if you get a value of y with a certain x, could you get an x with a value of y. I think the answer is YES....

RhondaSommer (rhondasommer):

no. i dont. im sorry. :/ i need to go eat supper too... @Zale101 @phi can you see if you can help this person please?

RhondaSommer (rhondasommer):

i told you they get the same answer when reversed. your confusing

OpenStudy (anonymous):

But when i plug in x=44 for y=(b*x-a)/(1+c*x), I get 12.96

OpenStudy (anonymous):

When i plug in x=55 in y=(b*x-a)/(1+c*x), I get 33... for y

OpenStudy (anonymous):

a= -3,516995576918316 b= 0,18734459582847 c= -0,0144436993225714